OpenStudy (yrivers36):

Use elimination to solve the system. 6x-4y=12 3x+5y=-6

OpenStudy (yrivers36):

I am not understanding how you obtained the answers

OpenStudy (anonymous):

Multiply equation (2) by \(-2\) and add it to equation (1), the system will become: \(6x-4y=12\) ... (1) \(-6x-10y=12\) .. (2) ----------------------- \(-14y=24 \implies y=-\frac{12}{7}\). Now plug this value of \(y\) in either equation to get the value of \(x\). Plugging in (1) yields \(6x-4(-\frac{12}{7})=12 \implies x=\frac{6}{7}\).

OpenStudy (anonymous):

So, \(y=-\frac{12}{7}\) and \(x=\frac{6}{7}\).

OpenStudy (yrivers36):

Thank you so much! I had the -12/7 but couldn't figure out rest of answer when I plugged it in. I kept getting 108/7

OpenStudy (anonymous):

I will do this part step by step, so you can see where you made your mistake. First, multiply both sides by \(7\) for simplicity: \(42x-4(-12)=84 \implies 42x+48=84 \implies 42x=36 \implies x=\frac{36}{42}\) Divide top and bottom by \(6\) to get \(x=\frac{6}{7}\).

OpenStudy (yrivers36):

Got it thank you

OpenStudy (anonymous):

You're welcome!